OpenStudy (anonymous):

Prove that the following limit exists and evaluate: lim (as n goes to infinity) ((n^2+3n+5)/(n^2+4n+7))^n

OpenStudy (amistre64):

compare leading terms

OpenStudy (amistre64):

just an idea

OpenStudy (amistre64):

n^2 ----- n^(2n) the bottom seems to be bigger than the top to me, what would this suggest?

OpenStudy (amistre64):

hmm, is that all to the ^n or just the bottom?

OpenStudy (anonymous):

it's the whole thing

OpenStudy (anonymous):

so it would be ((n^2/n^2))^n

OpenStudy (amistre64):

the ^n part messes with us, things change at different rates

OpenStudy (amistre64):

the wolf gives us 1/e but how to process it

OpenStudy (anonymous):

I have no idea.

OpenStudy (amistre64):

1 ---------- n^2+4n+7 | n^2+3n+5 -(n^2+4n+7) ------------ -2n-2 soo \[(1-2(\frac{n+1}{n^2+4n+7})^n\]

OpenStudy (amistre64):

so, this is prolly going to be related to: (1+f(n))^n

OpenStudy (amistre64):

\[(1+f(n))^n=\binom{n}{0}f^0(n)+\binom{n}{1}f^1(n)+...+\binom{n}{n}f^n(n)\]

OpenStudy (amistre64):

have you come across a thrm yet that says: (1+r/n)^(nt) as n to inf is equal to e^{rt} ?

OpenStudy (anonymous):

yes I just came across it.

OpenStudy (amistre64):

the less compact form of the binomial might be more useful here, just some ideas that are rolling around in my head

OpenStudy (anonymous):

i still don't see how that theorem gets 1/e like wolfram alpha does

OpenStudy (amistre64):

\[\binom{n}{k}=nCk=\frac{nPk}{k!}=\frac{n(n-1)(n-2)...(n-k+1)}{k(k-1)(k-2)...3.2.1}\] ugh, i used to know the process involved .... been along time

OpenStudy (anonymous):

would it be worth getting wolfram alpha pro so I could see the process?

OpenStudy (amistre64):

not to me, but thats your call

OpenStudy (amistre64):

can we make a taylor poly for this?

OpenStudy (anonymous):

......I don't know what that is

OpenStudy (amistre64):

its when we equate a polynomial to a function by taking derivatives to construct it

OpenStudy (anonymous):

I think we learn that in our differential equations class and we haven't covered it yet so I doubt that he would give a question where we had to solve it with that.

OpenStudy (amistre64):

im sure it deals with a compound interest limiting to a continuous interest setup

OpenStudy (anonymous):

I'm so confused

OpenStudy (amistre64):

1 + n f^0 + n(n-1)/2! f(n) + n(n-1)(n-2)/3! f^2(n) + ... \[(1+2\frac{-n-1}{n^2+4n+7})^n\] \[1+2n\left(\frac{-n-1}{n^2+4n+7}\right)+\frac{1}{2!}2n(n-1)\left(\frac{-n-1}{n^2+4n+7}\right)^2 \\ ~~~~~+\frac{1}{3!}2n(n-1)(n-2)\left(\frac{-n-1}{n^2+4n+7}\right)^3+...\] so im thinking as n to infinity we deal with the leading terms rules having a hard time with this new keyboard, might be better on paper

OpenStudy (amistre64):

\[1+\frac{(-2)^1}{1!}\frac{n(n)^1}{(n^2)^1}+\frac{(-2)^2}{2!}\frac{n^2(n)^2}{(n^2)^2}+\frac{(-2)^2}{3!}\frac{n^3(n)^3}{(n^2)^3}\] the n terms are all the same degrees so they go to 1 and we are left with \[\sum_{n=0}^{\infty}\frac{(-2)^n}{n!}\]

OpenStudy (amistre64):

almost .... that gets us 1/e^2 not 1/e

OpenStudy (zarkon):

\[\lim_{n\to\infty}\left(\dfrac{n^2+3n+5}{n^2+4n+7}\right)^n=\exp\left(\lim_{n\to\infty}n\ln\left(\dfrac{n^2+3n+5}{n^2+4n+7}\right)\right)\] \[=\exp\left(\lim_{n\to\infty}\dfrac{\ln\left(\dfrac{n^2+3n+5}{n^2+4n+7}\right)}{\dfrac{1}{n}}\right)\] now use L'Hospitals rule on the inside part

OpenStudy (anonymous):

\[\begin{align*}\lim_{n\to\infty}\left(\frac{n^2+3n+5}{n^2+4n+7}\right)^n&=\exp\left\{\ln\left[\lim_{n\to\infty}\left(\frac{n^2+3n+5}{n^2+4n+7}\right)^n\right]\right\} \end{align*}\] where (if you're unfamiliar with the notation, \(\exp[f(x)]=e^{f(x)}\)). Due to the continuity of the logarithmic function for \(x>0\), we can use the property, \[\lim_{x\to c}f(g(x))=f\left(\lim_{x\to c}g(x)\right)\] which in this case would be to say \[\begin{align*}\lim_{n\to\infty}\left(\frac{n^2+3n+5}{n^2+4n+7}\right)^n&=\exp\left\{\lim_{n\to\infty}\left[\ln\left(\frac{n^2+3n+5}{n^2+4n+7}\right)^n\right]\right\} \end{align*}\] (From here on I'll abbreviate the left side to \(L\).) Apply the exponent property for logarithms, i.e. \(\ln a^b=b\ln a\), giving \[\begin{align*}L&=\exp\left\{\lim_{n\to\infty}\left[n\ln\frac{n^2+3n+5}{n^2+4n+7}\right]\right\} \end{align*}\] Evaluating directly yields the indeterminate form \(0\times\infty\), so rearrange to a more useful expression: \[\begin{align*}L&=\exp\left\{\lim_{n\to\infty}\frac{\ln\dfrac{n^2+3n+5}{n^2+4n+7}}{\dfrac{1}{n}}\right\} \end{align*}\] Now evaluating directly yields the preferred \(\dfrac{0}{0}\) indeterminate form. Apply L'Hopital's rule. \[\begin{align*}L&=\exp\left\{\lim_{n\to\infty}\frac{\dfrac{\dfrac{(n^2+4n+7)(2n+3)-(n^2+3n+5)(2n+4)}{(n^2+4n+7)^2}}{\dfrac{n^2+3n+5}{n^2+4n+7}}}{-\dfrac{1}{n^2}}\right\}\\\\ &=\exp\left\{\lim_{n\to\infty}\dfrac{-n^2(n^2+4n+1)}{(n^2+3n+5)(n^2+4n+7)}\right\} \end{align*}\] Now you can consider the ratio of the coefficients of the leading terms.

OpenStudy (anonymous):

Note that this doesn't PROVE the limit exists, but it's just the computation of it. For the proof, it might be sufficient to show that the sequence converges (i.e. is bounded and monotonic).

OpenStudy (anonymous):

Seeing as we know the limit to be \(\dfrac{1}{e}\), we can make use of the limit definition of \(e^{-1}\) and draw some sort of conclusion from a comparison. \[e=\lim_{n\to\infty}\left(1+\frac{1}{n}\right)^n\]